OpenStudy (anonymous):

A square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 32 square feet. find the lengths of the sides of the poster and the sign.

OpenStudy (anonymous):

\[x^2-(x-2)^2=32\] solve for x

OpenStudy (anonymous):

careful with distributive law \[x^2-(x^2-4x+4)=32\] \[4x-4=32\] \[4x=36\] \[x=9\]

OpenStudy (anonymous):

x^2 - (x^2 - 4x + 4) = 35 4x - 4 = 35 4x = 36 ==> x = 9